OpenStudy (anonymous):

The height h in feet of a ball t seconds after being tossed upwards is given by the function h(t)=84t-16t^2. A. After how many seconds will it hit the ground? B. What is its maximum height? The topic of this problem is WORD PROBLEM INVOLVING QUADRATIC FUNCTION

OpenStudy (anonymous):

@Loser66 help please?

OpenStudy (anonymous):

@akitav

OpenStudy (anonymous):

@Nurali

OpenStudy (anonymous):

what would be the height of the ball when it hits the ground?

OpenStudy (anonymous):

110.25? i'm not pretty sure about my answer @akitav

OpenStudy (anonymous):

when the ball is on the ground height = 0, so the function which gives the height at any time would give h(t) = 0.\[84t-16t^2= 0\]taking t out \[t(84-16t) =0 \]so either t = 0 or (84-16t) = 0 which gives us t = 84/16 = 5.25 seconds (any unit for time given with the question?)

OpenStudy (anonymous):

Hmm. I see, isn't 2.63 is the seconds? Am i wrong? :3

OpenStudy (anonymous):

the function is \[h(t) = 84t-16t^2\] right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

then the answer for time when height = 0 or is t = 0 ( at the time of launch) or t = 5.25 seconds.

OpenStudy (anonymous):

Ohh. I understand now :) then after you get the seconds? I will substitute the t to the equation h(t)=84t-16t? Am i right or wrong?

OpenStudy (anonymous):

correct, for the second part. now one part to think about is the way to get the time for the highest point. it would spend exactly half of its time going up and half coming down. Since it took 5.25 sec in going up and coming down.\[5.25 = 2 \times t'\]solve for t' and put that value for answer to part b).

OpenStudy (anonymous):

What formula can i use for part b? Any hint please?

OpenStudy (anonymous):

Never mind for the hint :) hihi. I already know. ^_^ thank yooou so much @akitav :)